OpenStudy (anonymous):

The area of a rectangle is given by A = x2 + 4x – 5. Find the value of x if the total area of the rectangle is equal to 7 square units. a) x = 7 b) x = 2 c) x = 5 d) x = 1

OpenStudy (anonymous):

\(x^2+4x-5=7 => x^2+4x-12=0\) Then solve for \(x\)

OpenStudy (anonymous):

lets equate the equation to 7, since the area is equal to 7. x^2 + 4x -5 = 7 x^2 + 4x -5 -7 = 0 x^2 + 4x -12 = 0 Now since this is a 2 degree polynomial equation, you have to solve for x using the quadratic equation x = 2 and x = -6, since x cant be negative its x = 2;

OpenStudy (anonymous):

Thanks AS alway @javawarrior, & @Omniscience

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

quadratic formula* and your welcome :)